A business firms sells a good at the price of Rs 450.The firm has decided to reduce the
price of good to Rs 350.Consequently, the quantity demanded for the good rose from
25,000 units to 35,000 units. Calculate the price elasticity of demand.
1
Expert's answer
2020-11-03T11:17:06-0500

Answer: Price elasticity of demand is the measurement of a change in goods demanded with the price change. If the change is far from the original point, it is referred to as being elastic; simultaneously, if there is little change in the relationship between purchases and price change, we refer to it as being inelastic.

For this problem, I will use arc elasticity which is similar to the simpler price elasticity of demand (PED) but adds in the index problem. Arc elasticity uses a midpoint between the two points of a curve to measure the elasticity of a commodity.

To calculate PED, we first calculate the two midpoints;

Midpoint Quantity(Q)=(Q1+Q2)/2

Midpoint Price(P)=(P1+P2)/2

Therefore, Q=(25000+35000)/2=30000

P=(450+350)/2=400

Next,we calculate PED using the ordinary formula.

PED=%Change in Quantity(Q2-Q1)/Midpoint Q divided by %change in Price(P2-P1)/Midpoint P

=(35000-25000)/30000 divided by (350-450)/400

=0.33/-0.25

=-1.32